Please help me answer this question!
You have a bag containing 3 green marbles, 4 red marbles, and 2 yellow marbles. You select 1 marble randomly. What are the odds against selecting a green or yellow marble?
I tried adding them up and they came to 9. Are the odds against selecting a green or yellow marble mean that you're not going to pick them?
Thank you so much!

Answer by sudhanshu_kmr(1152)


probability that selecting a green or yellow P(A) = 5/9
P(A') = 1-5/9 = 4/9

odds against selecting a green or yellow = P(A')/P(A) = 4/5

if any doubt, you are welcome to contact.